OpenStudy (anonymous):

The partial sum of the first 6 terms of the arithmetic sequence with a1 = -6 and d = 4 is 24. True or False

Nnesha (nnesha):

do you know the formula to find other 5 terms of the sequence ?

OpenStudy (anonymous):

i think i have it but im not sure how to type it here haha

Nnesha (nnesha):

oh okay \[\huge\rm a_n = a_1+(n-1)d\] d= common difference a_1=first term n= number of term so to find 2nd term substitute n for 2

OpenStudy (anonymous):

so i just plug the a1 and d into that equation and add them, correct?

Nnesha (nnesha):

a_1 = -6 d=4 so plug them into the equation solve the parentheses first and then multiply by 4(d) and then add with -6

Nnesha (nnesha):

like for 2nd term \[\huge\rm a_2 = -6+(\color{ReD}{2}-1)4\] solve the parentheses first 2-1=1 -6+1(4) -6+4 =-2 so 2nd term is -2 find other 4 and then combine them

OpenStudy (anonymous):

i got the answer, thank you for your help again!